40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 30 00:58:19 2011, Koji, Summary, LSC, LSC whitening filter test chans_24_31_WeirdPhase.pdfOctopus.jpgTest_Inputs_Plugged_In.jpgContec_Tester_Board.jpg
    Reply  Thu Jul 7 02:23:59 2011, Jenne, Summary, LSC, LSC Whitening Filters have been fit AS55Q.png
       Reply  Thu Jul 7 15:34:44 2011, Jenne, Summary, LSC, LSC Whitening Filters have been fit 
Message ID: 4955     Entry time: Thu Jul 7 15:34:44 2011     In reply to: 4951
Author: Jenne 
Type: Summary 
Category: LSC 
Subject: LSC Whitening Filters have been fit 

Quote:

During the daytime either tomorrow or Friday I'll adjust the actual dewhitening filters to match the measured zpk values.

 I made a handy-dandy table showing the zpk values for each whitening filter in the wiki: New whitening filter page

Next on the whitening filter to-do list: actually put these values into the dewhitening filters in foton.

ELOG V3.1.3-